MississippiMan21:

The recipe for a batch of applesauce uses 2 and one-half pounds of apples. If Michael wants to make Three-fourths of a batch of applesauce, how many pounds of apples will he need? 3/10 of a pound 8/15 of a pound 1 7/8 of a pound 3 1/3 of a pound

Mercury:

set up a proportion \[\frac{ 2\frac{ 1 }{ 2 }~lbs }{ 1~batch }=\frac{ x~lbs }{ \frac{ 3 }{ 4 }~batch }\] and cross multiply to solve. the fractions make this a little more complicated, but you can convert 2 and 1/2 to an improper fraction, then multiply by 3/4